OpenStudy (anonymous):

Evaluate the following limit. lim sqrt(x+y) -3 (x,y) ->(3,6) /(x+y-9)

OpenStudy (anonymous):

reformatted: \[\lim_{(x,y) \rightarrow (3,6)}\frac{ \sqrt{x+y} -3}{ x+y-9 }\]

OpenStudy (dumbcow):

multiply by conjugate \[\frac{\sqrt{x +y} -3}{x+y-9}*\frac{\sqrt{x +y} +3}{\sqrt{x +y} +3}\]

OpenStudy (dumbcow):

the "x+y-9" terms will cancel out

OpenStudy (anonymous):

Thanks, lemme give this a try

OpenStudy (anonymous):

Thank you